# Active solids: defect self-propulsion without flow

### v1

@@ Abstract
Discussion. In active nematics $+1/2$, defects are generically expected to self-propel due to their polar symmetry. Such motility is transient in equilibrium, but becomes sustained in active systems. It has been studied extensively in fluids where defects are advected by the active flows they generate. Here we have proposed new mechanisms for self-sustained defect propulsion in active nematic solids where defects move relative to the elastic medium via local remodeling of the nematic texture.
Our work opens up many directions of future investigations. Additional elasto-nematic couplings that make the bend and splay Frank elastic constants differentially dependent on isotropic strain could mediate defect propulsion due to the bend-splay gradient across the $+1/2$ defect. This is in fact the more likely scenario for defect propulsion in Hydra, where no melting of order ahead of $+1/2$ defects has yet been observed.
It would also be interesting to examine the effect of noise, which we expect will lead to spontaneous defect unbinding analogously to the active fluid case [34, 36]. In biological contexts defect creation and unbinding might also result from coupling to morphogen concentration fields [37]. An important direction for future work is therefore to augment the current model by including such morphogen concentration fields which can, in turn, be sourced by the concentrated stresses near topological defects [38, 2]. Such a feedback loop may provide an alternate mechanism for the stabilization of the $+1$ defect at the head of Hydra [37, 2].
@@
Finally, self-propulsion through local melting may provide an explanation for a puzzling observation in cell monolayers, where defect motion often suggests extensile activity despite contractile forces at the cell scale [7, 6, 39]. Several explanations have been proposed [40, 41, 42, 43], including most recently that cell motion is primarily driven by propulsive traction with the substrate and that the nematic texture of cell shapes and associated topological defects is a passive response [44, 45]. The mechanism proposed here provides an alternative explanation, that does not require cell motility. The telltale sign of this mechanism is that defects move relative to the cells, as it is observed in regenerating Hydra  [1]. It will be interesting to investigate this in other tissues via the simultaneous use of defect tracking and Particle Image Velocimetry.
@@
We thank Kinneret Keren, Ananyo Maitra, Michael Moshe, and Marko Popović for insightful discussions and the Kavli Institute for Theoretical Physics for hospitality during the final stages of completion of this work. FB acknowledges support by the Gordon and Betty Moore Foundation post-doctoral fellowship (grant #2919). MCM was supported by the National Science Foundation award DMR-2041459.
@@ Appendix B Variational ansatz for defect propulsion
We use a variational formulation to find the dynamics of a $+1/2$ defect due to elasto-nematic coupling. To this end, we introduce the Rayleigh dissipation functional
@@
Physically, this equation expresses the power balance between the change of free energy and dissipation. To find the approximate equation of motion for the $+1/2$ defect, we substitute the ansatz $\mathbf{Q}(\mathbf{r},t)\approx\mathbf{Q}^{+\frac{1}{2}}[\mathbf{r}-\mathbf{r}_{0}(t)]$ into the variational equation, to obtain
@@ Appendix C Strain-alignment mediated defect propulsion
This result agrees well with numerics; see Fig. 4(b).
@@
showing the “duality” of the two defect propulsion mechanisms: the first operating through the combination of deviatoric nematogen-induced stress $(\sim\tilde{\alpha})$ and strain ordering $(\sim\chi_{B})$; the second operating through the combination isotropic nematogen-induced stress $(\sim\tilde{\alpha}_{B})$ and strain alignment $(\sim\chi)$.
@@
The isotropic strain concentrated at the defect core is sufficiently strong to completely “melt” the nematic order there via isotropic elasto-nematic coupling. In this case it is energetically favorable to localize the entire topological charge in the melted region. We can therefore estimate the active stress required to stabilize a $+1$ defect by requiring that the core is in the isotropic regime, i.e. $\chi_{B}u_{kk}(r=\sqrt{K})<-1$.
There is a family of $+1$ defects parametrized by the “phase” $\psi_{0}$
@@
where we assume that isotropic strain (and hence pressure) vanish at the boundary. Substituting into $\chi_{B}u_{kk}(r=\sqrt{K})<-1$ gives the estimated stability condition for $+1$ defects
@@
Comparing to numerics shows that Eq. (26) slightly underestimates the magnitude of $\tilde{\alpha}$ required to stabilize the $+1$ defect. Nonetheless, Eq. (26), provides several qualitative insights. First, in the region where the texture spontaneously twists into a spiral $+1$ defects can only be stabilized when the spiral pitch angle $\psi_{0}$ is sufficiently small. Indeed, there is only a very small regime with stable $+1$ spirals near the onset of spiral formation [see (ii) in Fig. 3]. Second, a $+1$ defect stabilizes more readily for a larger domain, since more active stress builds up over a larger area. Thus, the defect stabilization through elasto-nematic coupling is very different from stabilization through spatial confinement alone, where the latter requires a small domain size on the order of the nematic coherence length $\sqrt{K}$. In the presence of spring-like attachment to a substrate, elasticity becomes screened such that the system size in Eq. (26) is replaced by the screening length (see SI).

### v2

@@ Abstract
Discussion. In active nematics +1/2+1/2, defects are generically expected to self-propel due to their polar symmetry. Such motility is transient in equilibrium, but becomes sustained in active systems. It has been studied extensively in fluids where defects are advected by the active flows they generate. Here, we have proposed new mechanisms for self-sustained defect propulsion in active nematic solids where defects move relative to the elastic medium via local remodeling of the nematic texture.
@@
These mechanisms may provide an explanation for a puzzling observation in cell monolayers, where defect motion often suggests extensile activity despite contractile forces at the cell scale [9, 8, 45]. Several explanations have been proposed [46, 47, 48, 49], including most recently that cell motion is primarily driven by propulsive traction with the substrate and that the nematic texture of cell shapes and associated topological defects is a passive response [50, 51]. The mechanism proposed here provides an alternative explanation that does not require cell motility. The telltale sign of this mechanism is that defects move relative to the cells, as it is observed in regenerating Hydra [1]. It will be interesting to investigate this in other tissues via the simultaneous tracking of topological defects and individual cells.
In tissues, coupling between stress and nematic order might be mediated by intermediary biochemical processes and signaling molecules (morphogens). Coupling to morphogen concentration gradients provides additional mechanisms for defect unbinding and propulsion [52]. An important direction for future work is therefore to augment the current model by including such morphogen concentration fields which can, in turn, be sourced by the stress focused near topological defects [53, 2]. Such a feedback loop may provide an alternate mechanism for the stabilization of the +1+1 defect at the head of Hydra [52, 2].
Finally, in-plane stress can also drive out-of-plane deformations, generating 3D shapes [14], as has recently been utilized for tissue engineering [54]. Investigating the interplay of dynamic topological defects and 3D shape change is an important future direction with applications from morphogenesis and tissue engineering to soft robotics, where this interplay might guide the design of self-regulated actuators [55], pumps [56] or locomotion mechanisms [57, 58, 59].
@@
We thank Kinneret Keren, Ananyo Maitra, Carl Modes, Michael Moshe, and Marko Popović for insightful discussions and the Kavli Institute for Theoretical Physics for hospitality during the final stages of completion of this work. FB acknowledges support by the Gordon and Betty Moore Foundation post-doctoral fellowship (grant #2919). MO acknowledges support from the University of California Santa Barbara College of Creative Studies’ Francesc Roig Summer Undergraduate Research Fund. MCM was supported by the National Science Foundation award DMR-2041459.
@@ Appendix B Variational ansatz for defect propulsion
We use a variational formulation to find the dynamics of a +1/2+1/2 defect due to elasto-nematic coupling. To this end, we introduce the Rayleigh dissipation functional
@@
Physically, this equation expresses the power balance between the change of free energy and dissipation. To find the approximate equation of motion for the +1/2+1/2 defect, we substitute the ansatz $\mathbf{Q}(\mathbf{r},t)\approx\mathbf{Q}^{+\frac{1}{2}}[\mathbf{r}-\mathbf{r}_{0}(t)]$ into the variational equation, to obtain
@@ Appendix C Strain-alignment mediated defect propulsion
This result agrees well with numerics [ Fig. 4(b)].
@@
[…]
@@
The isotropic strain concentrated at the defect core is sufficiently strong to completely “melt” the nematic order there via isotropic elasto-nematic coupling. In this case it is energetically favorable to localize the entire topological charge in the melted region. We can therefore estimate the active stress required to stabilize a +1+1 defect by requiring that the core is in the isotropic regime, i.e. $\chi_{B}u_{kk}(r=\sqrt{K})<-1$.
There is a family of +1+1 defects parametrized by the “phase” $\psi_{0}$
@@
where we assume that isotropic strain (and hence pressure) vanish at the boundary. Substituting into $\chi_{B}u_{kk}(r=\sqrt{K})<-1$ gives the estimated stability condition for +1+1 defects
@@
Comparing to numerics shows that Eq. (26) slightly underestimates the magnitude of $\tilde{\alpha}$ required to stabilize the +1+1 defect. Nonetheless, Eq. (26), provides several qualitative insights. First, in the region where the texture spontaneously twists into a spiral +1+1 defects can only be stabilized when the spiral pitch angle $\psi_{0}$ is sufficiently small. Indeed, there is only a very small regime with stable +1+1 spirals near the onset of spiral formation [see (ii) in Fig. 3]. Second, a +1+1 defect stabilizes more readily for a larger domain, since more active stress builds up over a larger area. Thus, the defect stabilization through elasto-nematic coupling is very different from stabilization through spatial confinement alone, where the latter requires a small domain size on the order of the nematic coherence length $\sqrt{K}$. In the presence of spring-like attachment to a substrate, elasticity becomes screened such that the system size in Eq. (26) is replaced by the screening length (see SI).